twenty ounce of a 30% alcohol solution are mixed with 80 ounces of 20% alcohol solution. how much alcohol is in the mixture?
Answer by lwsshak3(11628)

twenty ounce of a 30% alcohol solution are mixed with 80 ounces of 20% alcohol solution. how much alcohol is in the mixture?
**
let x=% alcohol in the mixture
30%*20+20%*80=x*(20+80)
6+16=100x
x=22/100
x=22%
alcohol in the mixture=22%